What if the teacher asks what is the half of the sum of the brother of my cousin which is 12 and my sister which is 14 after 26 years from my birth. i am 8 years old.
Answer by ankor@dixie-net.com(22740)

What if the teacher asks what is the half of the sum of the brother of my cousin which is 12 and my sister which is 14 after 26 years from my birth.
i am 8 years old.
:
Find how many years from now,
26-8 = 18 yrs
:
.5[(12+18) + (14+18)] =
.5(30 + 32) =
.5(62) = 31 yrs